OpenStudy (anonymous):

help... will fan and medal!! Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1.Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

im not good with complete sentences

OpenStudy (amistre64):

ideally, this is some compounding rate: A = P(1+r)^t maybe even a continuous compounding rate: A = Pe^(rt)

OpenStudy (amistre64):

but without any data to go by, creating a function will be impossible

OpenStudy (dumbcow):

i think they just want you to make up a initial population and growth rate example: f(x) = 10(1.2)^x has initial value of 10 and growth rate of .2 or 20%

OpenStudy (anonymous):

then y does it say 1.2?? and not .2?

OpenStudy (dumbcow):

because to make a number increase by 20% you would multiply by 1.2 not .2 use the general form @amistre64 posted --> A = P(1+r)^x where "r" is growth rate

OpenStudy (anonymous):

ok thank u

OpenStudy (dumbcow):

yw